OpenStudy (studentchaitrali):

Trig.....need help!! If sinA-cosA=1 then, sin^3 A- cos^3 A is?

jimthompson5910 (jim_thompson5910):

\[\Large \sin(A) - \cos(A) = 1\] \[\Large (\sin(A) - \cos(A))^2 = 1^2\] \[\Large \sin^2(A) -2\sin(A)\cos(A) + \cos^2(A) = 1\] \[\Large \sin^2(A) + \cos^2(A) -2\sin(A)\cos(A) = 1\] \[\Large 1 -2\sin(A)\cos(A) = 1\] \[\Large -2\sin(A)\cos(A) = 0\] \[\Large \sin(A)\cos(A) = 0\]

jimthompson5910 (jim_thompson5910):

\[x^3 - y^3 = (x-y)*(x^2+xy+y^2)\] \[\sin^3(A) - \cos^3(A) = (\sin(A) - \cos(A))*(\sin^2(A) + \sin(A)\cos(A) + \cos^2(A))\] \[\sin^3(A) - \cos^3(A) = (\sin(A) - \cos(A))*(\sin^2(A)+\cos^2(A) + \sin(A)\cos(A))\] \[\sin^3(A) - \cos^3(A) = (\sin(A) - \cos(A))*(1 + \sin(A)\cos(A))\] \[\sin^3(A) - \cos^3(A) = (1)*(1 + 0)\] \[\sin^3(A) - \cos^3(A) = 1\]

OpenStudy (studentchaitrali):

Thnx